calculate the pH of 0.1 mol/L Sodium hydroxide(NaOH) with a Kb value= 36.


1
Expert's answer
2021-10-10T10:18:23-0400

"pH = \u2212log [H+]"

pH= - log (0.1)=1

pH= 14-1= 13